# From Many Models, One: Macroeconomic Forecasting with Reservoir Ensembles

### v1

@@ 5.5 Hyperparameter Sensitivity
The interesting takeaway from Figure 4 is that it suggests re-evaluating the $\alpha$ values used in Ballarin et al., 2024a . For all MFESN model classes, the most effective choices of the leak rate differ substantially from the original specifications. For instance, the baseline S-MFESNs models are constructed with $\alpha=0.1$ (see Table F.1), while panels (a)-(b) of Figure 4 suggest $\alpha\approx 0.7$ is a uniformly more robust choice. This insight, however, comes at a cost. Attempting to study multiple hyperparameters with a comparable level of precision, such as distinct monthly and daily frequency reservoir leak rates in M-MFESNs, would require the ensemble budget to grow exponentially. Model combination is thus potentially both more parsimonious – it does not require constructing a singular optimal model, but rather considering a fixed-budget set of models – and adaptive to potential shifts in individual model performance. This adaptivity is especially valuable for model classes like MFESNs, where randomness plays a considerable role in determining inductive model bias, which cannot be corrected ex-post through training. The same issue arises, for example, in neural networks with respect to architecture choices, activation functions, weight initialization, and regularization hyperparameters.
@@ 6 Conclusion
In this setting, our work makes significant contributions in both theoretical and empirical directions. First, we provide concentration bounds for stochastic ensemble losses that work for both i.i.d. and $\varphi$-mixing samples, and which exploit both Hoeffding-type and Bernstein-type results. Then, we apply these results to Follow-the-Leader and decreasing Hedge methods, for which we derive finite regret bounds. We then consider an application where model combination allows for robust and substantial performance improvements against previous state-of-the-art results: macroeconomic forecasting with mixed-frequency data using multi-frequency ESN models. We find that the ensemble approach is extremely effective at sharpening the performance of all models, with Follow-the-Leader and AdaHedge schemes being the overall most effective methods. Our empirical analysis also reveals that combinations can efficiently sidestep the question of hyperparameter tuning, which for (MF)ESN models is highly non-trivial.
@@ B.4.1 Combined Bounds for Hedge
Appendix C Direct Combination Methods
@@ Appendix D Algorithms
Appendix E Summary of Data
@@
Table E.1: Input and output variables, frequencies, and transformations (adapted from Ballarin et al., 2024a ).
Appendix F Summary of Models
@@
Table F.1: Models and ensembles applied in empirical forecasting exercises (for non-ensemble models see Table 4.1 in Ballarin et al., 2024a ). MFESN hyperparameters are defined for normalized state equations (11)-(12).
Appendix G Additional Tables
In this appendix, we provide results for the additional empirical experiments conducted under multistep forecasting scenarios. Tables G.1-G.3 report the relative MSE of the quarterly U.S. GDP growth predictions with respect to the sample mean for the exercise with the forecasting horizons taken as $h=2$, $h=4$, and $h=8$ quarters ahead, respectively. All competing models were implemented to produce the iterative $h$-steps-ahead forecasts (see Ballarin et al., 2024a for more details regarding the ESN-based models). It is important to emphasize that the hyperparameters for all the baseline MFESN models are tuned with cross-validation only for the prediction at $h=1$. This implies that baseline models are not expected to produce their best results with respect to the unconditional mean, AR(1), and DFM benchmarks. Our goal is to inspect the changes in the accuracy of the combinations of expert forecasts with respect to the performance offered by the standalone baseline models. We emphasize that in all combination schemes, the weights of the expert models in an ensemble are based on the past behavior in the $h$ steps ahead forecasting exercise. Our findings highlight that there are meaningful changes in relative improvements resulting from the use of combination strategies as the forecast horizon increases.
@@ Appendix G Additional Tables
Table G.1: Relative MSFE of quarterly U.S. GDP growth predictions with respect to the in-sample mean. Baselines are benchmarks and models in Ballarin et al., 2024a . Ensemble size is $K=1000$ for each MFESN specification. Performance changes in percentage with respect to baseline are shown in italic. Best performing model combinations are highlighted in bold.
@@
Table G.2: Relative MSFE of quarterly U.S. GDP growth predictions with respect to the in-sample mean. Baselines are benchmarks and models in Ballarin et al., 2024a . Ensemble size is $K=1000$ for each MFESN specification. Performance changes in percentage with respect to baseline are shown in italic. Best performing model combinations are highlighted in bold.
@@
Table G.3: Relative MSFE of quarterly U.S. GDP growth predictions with respect to the in-sample mean. Baselines are benchmarks and models in Ballarin et al., 2024a . Ensemble size is $K=1000$ for each MFESN specification. Performance changes in percentage with respect to baseline are shown in italic. Best performing model combinations are highlighted in bold.
Appendix H Additional Plots
(a) S-MFESN B – AdaHedge
(b) S-MFESN B – Follow-the-Leader
(c) M-MFESN B – AdaHedge
(d) M-MFESN B – Follow-the-Leader
[…]

### v2

@@ 5.4 Hyperparameter Sensitivity
These findings suggest revisiting the leak-rate choices in Ballarin et al., 2024a . Across MFESN classes, the most effective leak rates differ substantially from the original specifications. For instance, baseline S-MFESNs are constructed with $\alpha=0.1$ (Table I.1), while Figure 3(a)-(b) suggest the values around $0.7$ are a uniformly more robust choice. This insight, however, comes at a cost. Attempting to study multiple hyperparameters with a comparable level of precision, such as distinct monthly and daily frequency reservoir leak rates in M-MFESNs, would require the ensemble budget to grow exponentially. Model combination therefore offers a parsimonious and adaptive alternative, relying on a fixed-budget set of models rather than the construction of a single optimal specification. This adaptivity is particularly valuable for MFESNs, where randomness induces persistent model heterogeneity that cannot be corrected ex-post, a feature shared by many modern neural network architectures.
@@ 6 Conclusion
In this setting, our work makes contributions in both theoretical and empirical directions. First, we provide concentration bounds for stochastic ensemble losses under i.i.d. and $\varphi$-mixing assumptions, exploiting Hoeffding- and Bernstein-type results. Then, we apply these results to Follow-the-Leader and decreasing Hedge methods, for which we derive finite regret bounds. We then consider an application where model combination allows for robust and substantial performance improvements against previous state-of-the-art results: Macroeconomic forecasting with mixed-frequency data using multi-frequency ESN models. We find that the ensemble approach is extremely effective at sharpening the performance of all MFESN models, with Follow-the-Leader and AdaHedge schemes being the overall most effective schemes. Our empirical analysis also reveals that combinations can efficiently sidestep the hyperparameter tuning, which for (MF)ESN models is highly non-trivial.
@@ B.4.1 Combined Bounds for Hedge
Appendix C Doubling-trick Hedge
Specifically, let phase $r$ cover rounds $t\in[2^{r-1},2^{r}-1]$, and set
$\eta_{r}=\sqrt{\frac{8\log K}{S_{T}^{2}2^{r-1}}}.$
At the start of phase $r$, the forecaster resets the cumulative losses to zero and applies Hedge with constant rate $\eta_{r}$ until the end of the phase. Each phase therefore behaves as if the horizon were $2^{r-1}$ and the total number of phases up to round $T$ is at most $\lfloor\log_{2}T\rfloor+1$. Summing the regret bounds in (7) over all completed phases yields $\overline{R}_{T}\leq S_{T}\sqrt{2T\log K}+\mathcal{O}(S_{T}\log K)$, which matches the $\mathcal{O}(\sqrt{T\log K})$ worst-case rate of Hedge with known $T$.
Appendix D Direct Combination Methods
@@ Appendix E Algorithms
Appendix F Summary of Data
The dataset is collected from various sources, including the Federal Reserve Bank of St. Louis Monthly (FRED-MD) and Quarterly (FRED-QD) Macroeconomic Databases (see Stock and Watson,, 1996, 2002 for more details), as well as daily series sourced from Refinitiv Datastream. The selection of predictors in the medium-MD dataset includes indicators from ten macroeconomic and financial categories (see Stock and Watson,, 1996, 2002). The preprocessing of macroeconomic indicators follows the standard guidelines outlined in McCracken and Ng, (2016, 2020). The full data sample spans a period from January 1st, 1990, to December 31st, 2019. We concentrate on the evaluation up until the Great Financial Crisis, covering the period from 1990Q1 to 2007Q4. The testing and online learning sample spans 2008Q1-2019Q4, encompassing a total of 48 forecasting rounds.
@@ Appendix F Summary of Data
Table F.1: Input and output variables, frequencies, and transformations (adapted from Ballarin et al., 2024a ).
Appendix G Summary of Models
Model Name Description Specification
  AR(1) Autoregressive model of the output series estimated using OLS.
  DFM A Stock aggregation, VAR(1) factor process. Factors: 10
  DFM B Almon aggregation, VAR(1) factor process. Same as DFM A
[…]
EN-MFESN-RP Applied to all types of MFESN: 1000 distinct models generated with independently randomly drawn reservoir state coefficients. Same specification as the corresponding baseline MFESN model.
[…]
Table G.1: Models and ensembles applied in empirical forecasting exercises (for non-ensemble models see Table 4.1 in Ballarin et al., 2024a ). MFESN hyperparameters are defined for normalized state equations (11)-(12).
G.1 Combination Weights Dynamics
[…]
Our findings indicate that weight dynamics can differ significantly across different classes of MFESNs ensembles. In the upper row of Figure G.1, both AdaHedge and FTL weights for the S-MFESN A ensembles change substantially over the forecasting exercise. For AdaHedge (panel (a) in Figure G.1) the weight vector $\bm{\omega}_{\textnormal{{hdg}},t}$ remains very close to its equal-weight initialization until the second quarter of 2010. After that, a clearer ranking starts to emerge slowly, but ceases to stabilize. A similar pattern is visible for the Follow-the-Leader (panel (b) of Figure G.1), whose weights $\bm{\omega}_{\textnormal{{FTL}},t}$ fluctuate heavily until 2012Q2, after which the leader remains unchanged. By contrast, the dynamics of M-MFESN A ensembles’ weights are more stable. Panels (c) and (d) of Figure G.1 demonstrate that AdaHedge and FTL weights settle quickly and change slowly over time. For AdaHedge, the weights $\bm{\omega}_{\textnormal{{hdg}},t}$ stabilize after the end of the Great Financial Crisis downturn. Over 80% of the cumulative weight concentrates on just 5 experts, and the top-ranked expert receives over 30% of the weight as early as 2009Q1. Furthermore, the weights $\bm{\omega}_{\textnormal{{FTL}},t}$ of the FTL demonstrate quick convergence to a stable leader within less than one year from the start of the prediction exercise, signaling that the ensemble contains an M-MFESN A model which achieves consistent optimality in terms of MSFE.
It is important to emphasize that these marked differences in weight dynamics do not imply large differences in overall forecasting performance. Table 2 reports that MFESN A-type results for FTL and AdaHedge across single- and multi-reservoir architectures show much smaller MSFE gaps than, for example, across A- and B-type architectures. From an interpretability perspective, more stable weights provide clearer insight into the mechanism of the ensemble: Figure G.1(d) strongly suggests an online form of model selection carried out iteratively during the forecasting exercise. An open question is hence whether there exist inherent trade-offs between ensemble “interpretability” (e.g., provably simple weights dynamics for model instances from given classes of nonlinear models) and predictive performance. We defer the treatment of this question to future work.
(a) S-MFESN A – AdaHedge
(b) S-MFESN A – Follow-the-Leader
(c) M-MFESN A – AdaHedge
(d) M-MFESN A – Follow-the-Leader
[…]
(a) S-MFESN B – AdaHedge
(b) S-MFESN B – Follow-the-Leader
(c) M-MFESN B – AdaHedge
(d) M-MFESN B – Follow-the-Leader
[…]
Appendix H Additional Tables
In this appendix, we provide results for the additional empirical experiments conducted under multistep forecasting scenarios. Tables J.1-J.3 report the relative MSE of the quarterly U.S. GDP growth predictions with respect to the sample mean for the exercise with the forecasting horizons taken as $h=2$, $h=4$, and $h=8$ quarters ahead, respectively. All competing models were implemented to produce the iterative $h$-steps-ahead forecasts (see Ballarin et al., 2024a for more details regarding the ESN-based models). It is important to emphasize that the hyperparameters for all the baseline MFESN models are tuned with cross-validation only for the prediction at $h=1$. This implies that baseline models are not expected to produce their best results with respect to the unconditional mean, AR(1), and DFM benchmarks. Our goal is to inspect the changes in the accuracy of the combinations of expert forecasts with respect to the performance offered by the standalone baseline models. We emphasize that in all combination schemes, the weights of the expert models in an ensemble are based on the past behavior in the $h$ steps ahead forecasting exercise. Our findings highlight that there are meaningful changes in relative improvements resulting from the use of combination strategies as the forecast horizon increases.
More specifically, for $h=2$, similar to the case of $h=1$ in Table 2, the FTL combination strategy together with the AdaHedge algorithm again appear to dominate all other combination schemes for the corresponding baseline models, significantly improving the forecasting accuracy of the latter once used in an ensemble. Starting from $h=4$ onward, several differences are noticeable. First, the expert selected by FTL as the leader (that is, the one that produces the most accurate predictions measured in terms of the cumulative loss) ceases to be the best predictor in the subsequent period. While FTL can adapt quickly to changes in the best expert for one-step-ahead predictions, its ability to track the best expert over longer horizons diminishes. Second, the AdaHedge and DecHedge strategies maintain strong performance for longer prediction horizons, especially for the case of EN-MFESN-RP ensembles, where only the parameters of the state equation are resampled.
—  Horizon $h=2$  —
[…]
Table H.1: Relative MSFE of quarterly U.S. GDP growth predictions with respect to the in-sample mean. Baselines are benchmarks and models in Ballarin et al., 2024a . Ensemble size is $K=1000$ for each MFESN specification. Performance changes in percentage with respect to baseline are shown in italic. Best performing model combinations are highlighted in bold.
—  Horizon $h=4$  —
[…]
Table H.2: Relative MSFE of quarterly U.S. GDP growth predictions with respect to the in-sample mean. Baselines are benchmarks and models in Ballarin et al., 2024a . Ensemble size is $K=1000$ for each MFESN specification. Performance changes in percentage with respect to baseline are shown in italic. Best performing model combinations are highlighted in bold.
—  Horizon $h=8$  —
[…]
Table H.3: Relative MSFE of quarterly U.S. GDP growth predictions with respect to the in-sample mean. Baselines are benchmarks and models in Ballarin et al., 2024a . Ensemble size is $K=1000$ for each MFESN specification. Performance changes in percentage with respect to baseline are shown in italic. Best performing model combinations are highlighted in bold.
Start Date T Code Name Description
Quarterly
31/03/1959 5 GDPC1 Y Real Gross Domestic Produce
Monthly
30/01/1959 5 INDPRO XM1 Industrial Production Index
30/01/1959 5 PAYEMS XM4 Payroll All Employees: Total nonfarm
30/01/1959 4 HOUST XM5 Housing Starts: Total New Privately Owned
30/01/1959 5 RETAILx XM7 Retail and Food Services Sales
31/01/1973 5 TWEXMMTH XM11 Nominal effective exchange rate US
30/01/1959 2 FEDFUNDS XM12 Effective Federal Funds Rate
30/01/1959 1 BAAFFM XM14 Moody’s Baa Corporate Bond Minus FEDFUNDS
30/01/1959 1 COMPAPFFx XM15 3-Month Commercial Paper Minus FEDFUNDS
30/01/1959 2 CUMFNS XM2 Capacity Utilization: Manufacturing
30/01/1959 2 UNRATE XM3 Civilian Unemployment Rate
30/01/1959 5 AMDMNOx XM8 New Orders for Durable Goods
31/01/1978 2 UMCSENTx XM9 Consumer Sentiment Index
30/01/1959 6 WPSFD49207 XM10 PPI: Finished Goods
30/01/1959 1 AAAFFM XM13 Moody’s Aaa Corporate Bond Minus FEDFUNDS
30/01/1959 1 TB3SMFFM XM16 3-Month Treasury C Minus FEDFUNDS
30/01/1959 1 T10YFFM XM17 10-Year Treasury C Minus FEDFUNDS
30/01/1959 2 GS1 XM18 1-Year Treasury Rate
30/01/1959 2 GS10 XM19 10-Year Treasury Rate
Daily
30/01/1959 8 DJINDUS XD3 DJ Industrial price index
31/12/1963 8 S&PCOMP XD1 S&P500 price index
01/05/1982 1 ISPCS00-S&PCOMP* XD2 S&P500 basis spread
11/09/1989 8 SP5EIND XD4 S&P Industrial price index
31/12/1969 8 GSCITOT XD5 Spot commodity price index
10/01/1983 8 CRUDOIL XD6 Spot price oil
02/01/1979 8 GOLDHAR XD7 Spot price gold
30/03/1982 8 WHEATSF XD8 Spot price wheat
01/11/1983 8 COCOAIC,COCINU** XD9 Spot price cocoa
30/03/1983 1 NCLC.03-NCLC.01 XD10 Futures price oil term structure
30/10/1978 1 NGCC.03-NGCC.01 XD11 Futures price gold term structure
Table H.4: Input and output variables, frequencies, and transformations (adapted from Ballarin et al., 2024a ).
Appendix I Summary of Models
@@
Table I.1: Models and ensembles applied in empirical forecasting exercises (for non-ensemble models see Table 4.1 in Ballarin et al., 2024a ). MFESN hyperparameters are defined for normalized state equations (11)-(12).
Appendix J Additional Tables
In this appendix, we provide results for the additional empirical experiments conducted under multistep forecasting scenarios. Tables J.1-J.3 report the relative MSE of the quarterly U.S. GDP growth predictions with respect to the sample mean for the exercise with the forecasting horizons taken as $h=2$, $h=4$, and $h=8$ quarters ahead, respectively. All competing models were implemented to produce the iterative $h$-steps-ahead forecasts (see Ballarin et al., 2024a for more details regarding the ESN-based models). It is important to emphasize that the hyperparameters for all the baseline MFESN models are tuned with cross-validation only for the prediction at $h=1$. This implies that baseline models are not expected to produce their best results with respect to the unconditional mean, AR(1), and DFM benchmarks. Our goal is to inspect the changes in the accuracy of the combinations of expert forecasts with respect to the performance offered by the standalone baseline models. We emphasize that in all combination schemes, the weights of the expert models in an ensemble are based on the past behavior in the $h$ steps ahead forecasting exercise. Our findings highlight that there are meaningful changes in relative improvements resulting from the use of combination strategies as the forecast horizon increases.
@@
Table J.1: Relative MSFE of quarterly U.S. GDP growth predictions with respect to the in-sample mean. Baselines are benchmarks and models in Ballarin et al., 2024a . Ensemble size is $K=1000$ for each MFESN specification. Performance changes in percentage with respect to baseline are shown in italic. Best performing model combinations are highlighted in bold.
@@
Table J.2: Relative MSFE of quarterly U.S. GDP growth predictions with respect to the in-sample mean. Baselines are benchmarks and models in Ballarin et al., 2024a . Ensemble size is $K=1000$ for each MFESN specification. Performance changes in percentage with respect to baseline are shown in italic. Best performing model combinations are highlighted in bold.
@@
Table J.3: Relative MSFE of quarterly U.S. GDP growth predictions with respect to the in-sample mean. Baselines are benchmarks and models in Ballarin et al., 2024a . Ensemble size is $K=1000$ for each MFESN specification. Performance changes in percentage with respect to baseline are shown in italic. Best performing model combinations are highlighted in bold.